OpenStudy (anonymous):

How does the division rule for exponents help in understanding why anything to the zero power is 1?

OpenStudy (anonymous):

For the division rule you subtract the exponents. For the difference to be 0 the exponents have to be equal so that means the numerator is equal to the denominator and any time the numerator is equal to the denominator the fraction simplifies to 1.

OpenStudy (anonymous):

what rohangrr said. i math you could write \[1=\frac{b^n}{b^n}=b^{n-n}=b^0\]